OpenStudy (anonymous):

show that the sum of a complex number and its conjugate is twice the real part of the number, and the difference is twice the imaginary part of the number.

OpenStudy (anonymous):

do u knw what is complex no.?

OpenStudy (anonymous):

let me know that first?

OpenStudy (perl):

(a+bi) - (a-bi) = 2b i

OpenStudy (anonymous):

@manishsatywali : yes i know it..

OpenStudy (anonymous):

so fr the moment u have been provided that : for a complex no . A+ Bi

OpenStudy (anonymous):

what will be it's conjugate ???? tell me asap

OpenStudy (anonymous):

@perl : thanks..

OpenStudy (anonymous):

@manishsatywali : a-bi

OpenStudy (anonymous):

k now according to th equestion there r two conditions first diffrence and sum of complex no. and it's conjugate ......ok

OpenStudy (perl):

oh i messed up the first one

OpenStudy (perl):

(a+bi) + ( a-bi) = 2a (a+bi) - (a-bi) = 2b i

OpenStudy (anonymous):

frst let's focus on the sum of complex no. and it's conjugate

OpenStudy (perl):

manish, dont sweat it bro. this is done

OpenStudy (perl):

where is my medal?

OpenStudy (perl):

(a+bi) + ( a-bi) = 2a (a+bi) - (a-bi) = 2b i

OpenStudy (anonymous):

@perl pls...dn't give direct answer...............i getting whether she everything about the question or not right from base level

OpenStudy (anonymous):

@manishsatywali : i agree..

OpenStudy (perl):

manish, dont sweat it bro.

OpenStudy (perl):

manish, i got this bro

OpenStudy (anonymous):

nw sum (A-Bi ) + (A-Bi).........do it fast ?????????????@meejoy

OpenStudy (perl):

no thats wrong manish

OpenStudy (perl):

manish, here (a+bi) + ( a-bi) = 2a (a+bi) - (a-bi) = 2b i

OpenStudy (perl):

a = Re(z) , b = Im (z) ,

OpenStudy (anonymous):

@perl bro pls. w8

OpenStudy (anonymous):

@meejoy do the sum

OpenStudy (perl):

ok sorry :) message me if you have question

OpenStudy (anonymous):

@perl and @manishsatywali : this is for studying purpose... but it seems this just gave me a mess. >_<

OpenStudy (anonymous):

@meejoy.........pls do the sum

OpenStudy (anonymous):

@perl and @manishsatywali : stop the wild talk ..

OpenStudy (perl):

wild talk?

OpenStudy (anonymous):

2A-2Bi

OpenStudy (anonymous):

srry @meejoy u gt the wrong answer + Bi and - Bi must cancel

OpenStudy (anonymous):

you ask me to get the sum of this right ? (A-Bi ) + (A-Bi) .. how come that Bi must be cancel???

OpenStudy (anonymous):

@manishsatywali : are you still here?

OpenStudy (anonymous):

@perl : will you pls help me to know and understand the right answer and solution? thanks..

OpenStudy (anonymous):